# Topological Optimal Transport for Geometric Cycle Matching

### v1

@@ 5.2 Geometric cycle matching
In our case, the topological feature coupling $\pi^{e}$ yields a matching between homology classes which is informed by proximity of points creating the corresponding geometric cycles. Figure 6 shows comparisons between the Wasserstein matching on persistence diagrams (PD Wasserstein) and joint geometric-topological matching (TpOT). Similarly to Figure 4, the first examples concerns mapping classes between point clouds $X$ (Figure 6(A)) and $X^{\prime}$ (Figure 6(B)) having almost identical persistence diagrams (Figure 6(C)). In this case, joint matching of geometric and topological features by TpOT allows us to preserve spatial proximity between representative cycles, see Figure 6(C). This can be particularly important when homology classes in a persistence diagram are indistinguishable, as shown in the example in Figure 6(D-G). Here the TpOT matching (Figure 6(G)) correctly maintains spatial proximity between representative cycles, in contrast with PD Wasserstein matching (Figure 6(F)) which is essentially random.
Finally, we consider the problem of tracking topological features as a system evolves over time. Figure 7(A) shows a piecewise linear spatial curve forming a trefoil knot as it undergoes thermal relaxation. The simulation consists of $20$ steps and it was generated using KnotPlot [41]. At each time $t_{i}$, the knot takes a spatial configuration $C_{i}$, that can be represented as the point cloud $X_{i}$ of its vertices. At each step, we interpret $X_{i}$ as the topological network obtained by computing 1-dimensional PH and PH-hypergraph. We can then match homology classes between consecutive persistence diagrams $D_{i},D_{i+1}$ using PD Wasserstein distance and $d_{\rm TpOT}$, respectively, see Figure 7(B). The TpOT matching significantly improves the accuracy of generator matchings compared to PD Wasserstein, as shown in Figure 7(B). Since the ground truth correspondence of points in the curve is known, in Figure 7(C) we show the correlation between generators and their images under the PD Wasserstein as well as TpOT matchings. We find that the TpOT matching provides a clear improvement over the PD Wasserstein matchings, which appears to be close to random.
@@
Consider two point clouds, and the convex $GW$ geodesic connecting them (see [43]). Consider now the Vietoris-Rips filtration along the geodesic path. A natural question is whether the path induced by the persistent diagrams follows a convex geodesic. A counterexample to this question is given by Figure 3, as the GW-mapping between points suggests that the GW-geodesic “breaks” the loops, and thus, the four persistent homology classes. This can be verified by computing the path of Vietoris-Rips PDs obtained by inputting the various interpolation metrics. The same question applies for $TpOT$-geodesics. In this context, the answer depends on whether the $GW$-term contributes to cycle matching or not. When it doesn’t (e.g. Figure 3(E)), persistent diagrams computed from Vietoris-Rips filtration along the TpOT-geodesic do indeed follow the convex-geodesic in the persistent diagram space. When GW contributes (e.g. Figure 4 and Figure 6), then the persistent diagram path is non geodesic.
@@ 6 Conclusion
Inspired by recent advances in computational topology [2] and optimal transport for structured objects [48, 39, 11], we propose to encode geometric and topological information about a point cloud jointly in a hypergraph structure. Our approach lays the foundation for a topology-driven analysis of the geometry of gauged measure spaces. As in the co-optimal setting [39, 11], the TpOT problem outputs a pair of coupled transport plans. The first one provides a matching of the underlying metric spaces (point clouds) that preservers topological features. The second one matches the topological features (cycles representing persistent homology classes) in a way that maintains local metric structure. This geometric cycle matching represents a key novelty of our method, and, to the best of our knowledge, it is the first one taking into account the geometric nature of data, and the spatial interconnectivity of generating cycles.
Constructing PH-hypergraphs, and thus the TpOT pipeline, depends on computing explicit cycles representing homology classes. It is well known that, even in the case of simplicial complexes, generating cycles are far from being unique, and that different choices can potentially bring important biases in the subsequent analysis [29]. Given their potential and demonstrate usefulness, the search for new methods of producing “good” representatives, and the research into homology generators in general [5, 16, 19, 29, 34, 49, 2], is a very active sub-field of topological data analysis. New, important advancements in this area currently provide a wide variety of algorithms to compute representatives, that can all be implemented in the TpOT pipeline, depending on the specific goal or application. In addition to that, the experimentally shown robustness of the hyperTDA method [2] suggests that, when dealing with systems sufficiently complex from a topological point of view, the qualitative behaviour of a TpOT induced matching should be robust by changes to the representative cycles.

### v2

@@ 5.2 Geometric cycle matching
In our case, the topological feature coupling $\pi^{e}$ yields a matching between homology classes which is informed by the proximity of points creating the corresponding geometric cycles. We showcase this in Figure 6 where we compare the standard Wasserstein matching on the persistence diagram space (“PD Wasserstein”) and our TpOT matchings. In Figure 6(a-b) consider two point clouds $X$ and $X^{\prime}$, each consisting of three cycles of different sizes and arranged differently in space. Consequently, the persistence diagrams of $X$ and $X^{\prime}$ are nearly identical, despite the relative positioning of the cycles being different. Joint matching of geometric and topological features using TpOT produces a mapping between generators that preserves their layout. In contrast, Wasserstein matching on the persistence diagrams alone matches features by their persistence alone, breaking the spatial arrangement. As a more extreme example, in Figure 6(c) we consider two point clouds $X$, $X^{\prime}$ with 9 identical loops each. As before, TpOT recovers a matching of generators that preserves their relative ordering while Wasserstein matching on the persistence diagrams produces a shuffled ordering. In all of these examples, measure topological networks $P_{X},P_{X^{\prime}}$ are constructed using affinity matrices of squared Euclidean distances and binary hypernetwork incidence function. TpOT matchings are computed as before using $\varepsilon_{v}=3\times 10^{-3},\varepsilon_{e}=10^{-2},\alpha=0.5,\beta=1.0$.
Finally, we consider the problem of tracking topological features as a system evolves over time. Figure 7(a) shows a piecewise linear spatial curve forming a trefoil knot as it undergoes thermal relaxation. The simulation consists of $20$ steps generated using KnotPlot [52]. At each time $t_{i}$, the knot takes a spatial configuration $C_{i}$, that can be represented as the point cloud $X_{i}$ of its vertices. At each step, we construct from $X_{i}$ a measure topological network $P_{X_{i}}$ obtained by computing 1-dimensional PH. Similar to the example in Figure 5, we choose to use a Gaussian kernel for pairwise affinity matrices, and use the Laplacian-smoothed binary incidence matrices. We can then match homology classes between consecutive temporal instants $(t_{i},t_{i+1})$ by either Wasserstein matchings on persistence diagrams or TpOT. Since the ground truth correspondence of points in the curve is known, in Figure 7(b) we show the correlation between generators. We find that the TpOT matching significantly improves the accuracy of generator matchings compared to PD Wasserstein. Finally, in Figure 7(c) we show the images of significant generators under the mappings found by PD Wasserstein as well as TpOT. We find that the TpOT matching provides a clear improvement over the PD Wasserstein matching.
Reliably tracking topological features as a system evolves over time offers the key advantage of identifying structural features resilient to dynamic changes. For spatial curves, this capability has clear implications for studying the folding and dynamics of (bio)polymers. In a polymer with a specific—potentially knotted—spatial configuration, are regions that take longer to untangle biophysically significant, and if so, why? A natural setting where this question is particularly relevant is the case of proteins. Folding intermediates are difficult to obtain, both experimentally and via simulations, and similar topological considerations can inform and enhance the analysis of protein folding.
@@
Consider two point clouds, and the convex $GW$ geodesic connecting them (see [54]). Consider now the Vietoris-Rips filtration along the geodesic path. A natural question is whether the path induced by the persistent diagrams follows a convex geodesic. A counterexample to this question is given by Figure 3, as the GW-mapping between points suggests that the GW-geodesic “breaks” the loops, and thus, the four persistent homology classes. This can be verified by computing the path of Vietoris-Rips PDs obtained by inputting the various interpolation metrics. The same question applies for $TpOT$-geodesics. In this context, the answer depends on whether the $GW$-term contributes to cycle matching or not. When it doesn’t (e.g. Figure 3(E)), persistent diagrams computed from Vietoris-Rips filtration along the TpOT-geodesic do indeed follow the convex-geodesic in the persistent diagram space. When GW contributes (e.g. Figure 4 and Figure 6), then the persistent diagram path is non-geodesic.
5.3 Higher-degree homology
Note that the TpOT approach is not constrained to 1D PH; it can be applied more broadly to higher-dimensional PH data. The structure and complexity of PH-hypergraphs varies with the homology dimension [2], and for data lying in higher-dimensional spaces, the information from homology in higher degrees may provide more nuanced insights. As a concrete example, recent work by a subset of the authors and collaborators demonstrates that 2-dimensional cycles are effective in detecting damaging mutations in protein structures [40]. For clarity and conciseness, and because this setting is directly relevant to the experiments we present, we have focused on the 1D case.
5.4 Computational complexity
Our method computes an approximate solution of a non-convex quadratic program, which is NP-hard to solve exactly. Our iterative algorithm for finding an approximate solution relies on solving successive linear optimal transport subproblems, which are at least $O(n^{3}\log n)$ in the exact case. The entropy regularised case is slightly more efficient, see e.g. [46]. We remark that our method currently stores full cost matrices in memory and is thus $O(n^{2})$ in space; for large point clouds speedups could be used. Persistent homology computations are known to be computationally very expensive (see e.g. [56, 35]). As an example, the computation of first-degree homology requires analysing $O(n^{3})$ $2$-simplices. For these computations, we are using the Julia package Ripserer [16], one of the top performing softwares currently available.
5.5 Generating cycles
Constructing PH-hypergraphs, and thus the TpOT pipeline, depends on computing explicit cycles representing homology classes. It is well known that, even in the case of simplicial complexes, generating cycles are far from being unique, and that different choices can potentially bring important biases in the subsequent analysis [38]. Ideally, one would want to work with some minimal generating basis, or have some sort of preferred or best-practice method to choose cycle representatives. Unfortunately, in general, finding minimal generators is an NP-hard problem [18, 38]. In practice, we observe good results with the representatives found by the software Ripserer [16], a package which has enjoyed widespread adoption in the TDA community. In addition to that, the experimentally shown robustness of the hyperTDA method [2] suggests that, when dealing with systems sufficiently complex from a topological point of view, the qualitative behaviour of a TpOT induced matching should be robust by changes to the representative cycles. Further, the implementation of the approach by Grande and Schaub [31] assures an increased robustness of the overall pipeline.
Given their potential and demonstrated usefulness, the search for new methods of producing “good” representatives, and the research into homology generators in general [6, 18, 22, 38, 44, 60, 2], is a very active sub-field of topological data analysis. New, important advancements in this area currently provide a wide variety of algorithms to compute representatives, that can all be implemented in the TpOT pipeline, depending on the specific goal or application.
@@ 6 Conclusion
Inspired by recent advances in computational topology [2] and optimal transport for structured objects [58, 49, 13], we propose to encode geometric and topological information about a point cloud jointly in a hypergraph structure. Our approach lays the foundation for a topology-driven analysis of the geometry of gauged measure spaces. As in the co-optimal setting [49, 13], the TpOT problem outputs a pair of coupled transport plans. The first one provides a matching of the underlying metric spaces (point clouds) that preserves topological features. The second one matches the topological features (cycles representing persistent homology classes) in a way that maintains local metric structure. This geometric cycle matching represents a key novelty of our method, and, to the best of our knowledge, it is the first one taking into account the geometric nature of data, and the spatial interconnectivity of generating cycles.